OpenStudy (firejay5):

A college with a graduating class of 3,600 students in the year 2012, predicts that it will have a graduating class of 4,862 in 2016. 1. Write an exponential function to model the # of students y in the graduating class t years after 2012. 2. What is the growth factor of this college?

OpenStudy (rational):

what have you tried so far

OpenStudy (firejay5):

@rational I have tried nothing, because I don't know how to use exponential function in a world application problem

OpenStudy (rational):

I'm pretty sure there mus tbe something about this in your notes/textbook and if you see lesson pages you might actually find the answer

OpenStudy (firejay5):

is 1. 4862= 3600(something)^x ; y = 3600(1.35)^x ----> @rational

OpenStudy (firejay5):

@rational what's #2?

OpenStudy (firejay5):

Is it growing 1.35 students a year

OpenStudy (rational):

\[\large y = AB^{t}\]

OpenStudy (rational):

starting population = 3600 so A = 3600 \[\large y = 3600B^{t}\]

OpenStudy (firejay5):

I need #2

OpenStudy (rational):

your part a is wrong

OpenStudy (rational):

you need part a to figure out part b

OpenStudy (firejay5):

what is wrong with part a? I got to get ready for school, so could you please give me the answer

OpenStudy (rational):

the prediction for population in `four` years is `4862` so you need to plugin t = 4 and y = 4862 in above equation and solve B : \[4862 = 3600B^4\] solve \(B\)

OpenStudy (firejay5):

1.35x^4 right

OpenStudy (rational):

you will get B = 1.078

OpenStudy (firejay5):

then take the fourth root to both sides and you get 1.08

OpenStudy (rational):

Yes thats it!

OpenStudy (firejay5):

would it be y = 3600 (1.08)^x

OpenStudy (firejay5):

I need b's answer like now

OpenStudy (rational):

so the answer for part 1 would be \[\large y = 3600 (1.078)^t\]

OpenStudy (rational):

for second part, the growith factor is simply B : 1.078

OpenStudy (firejay5):

1.078 what?

OpenStudy (rational):

yes thats the factor, growth factor is a number

OpenStudy (firejay5):

okay thank you helped me so much! :D You're amazing

OpenStudy (rational):

the population grows by 7.8% each year

OpenStudy (rational):

yw